OpenStudy (anonymous):

How do you evaluate this integral using substitution ? ∫x/x+3dx

OpenStudy (turingtest):

\[u=x+3\implies x=?\]

OpenStudy (turingtest):

it u=x+3 what does x equal ???

OpenStudy (anonymous):

x= u-3 ?

OpenStudy (anonymous):

dont go for substitution...do it like this......... (x+3)-3/(x+3)= (x+3) /(x+3) - 3/(x+3) which inturn equals to 1- 3/(x+3)......and integration of 1 will be x.....and integration of 3/x+3 will be 3 ln(x+3)

OpenStudy (anonymous):

By Substitution: \[ t = x+3\] \[ x = t - 3 \] \[ dt = dx \] Substitute them in the original equation: \[ \int \frac{x}{x+3} dx \] \[ = \int \frac{t-3}{t} dt \] \[ = \int (1 - \frac{3}{t})dt \] \[ = \int dt - \int \frac{3}{t} dt \] \[ = t - 3ln(t) +c_1 \] \[ = x + 3 - 3ln(x+3) + c_1 \] \[ = x - 3ln(x+3) + c \] Simple Solving: \[ \int \frac{x}{x+3} dx \] \[ = \int \frac{x+3-3}{x+3} dx \] \[ = \int (1 - \frac{3}{x+3}) dx\] \[ = \int dx - \int \frac{3}{x+3} dx \] \[ = x - 3 ln(x+3) + c \]